Cyberdyne Systems Corporation uses a standard-costing system. The production budget for the fiscal year (FY) ended 30 November 2018 was based on 200,000 units of output. Each unit requires two standard hours of manufacturing labor for completion. Total overhead was budgeted at $900,000 for the year, and the budgeted fixed overhead rate was $3 per unit of output. Both fixed and variable overhead are allocated to the product on the basis of direct manufacturing labor-hours. The actual data for the fiscal year ended 30 November 2018 are as follows: actual production in units, 198,000; actual direct manufacturing labor-hours, 440,000; actual variable overhead, $352,000; actual fixed overhead, $575,000.

The Cyberdyne fixed overhead spending variance for FY2018 is closest to:

a. $19,000 F

b. $19,000 U

c. $25,000 F

d. $25,000 U

The Cyberdyne output level variance for FY2018 is closest to:

a. $25,000 F

b. $55,000 U

c. $19,000 F

d. $6,000 U

The Cyberdyne variable overhead spending variance for FY2018 is closest to:

a. $20,000 U

b. $20,000 F

c. $22,000 U

d. $22,000 F

The Cyberdyne variable overhead efficiency variance for FY2018 is closest to:

a. $33,000 U

b. $35,520 F

c. $66,000 U

d. $35,200 U
Answer #1

Fixed overhead spending variance  

(C) $25,000 F

Fixed Overhead spending variance = Actual overhead - Budgeted overhead
                                                                        = 575,000 - 600,000
                                                                        = (25,000 ) F
Budgeted overhead = Budgeted units * Budgeted overhead rate
                                          =200,0000 * 3
                                         = 600,000

Output level Variance

(D) $6,000 U

Output volume variance = (Actual output - Standard output ) * Budgeted overhead rate
                                                  = ( 98,000 - 100,000) * 3
                                                 =   6,000 U

Variable overhead spending variance

(D) 22,000

Variable overhead spending variance = Actual Hours worked * ( Actual overhead rate- standard overhead rate)
                                                                             = 440000 * ( 0.80 -0.75)
                                                                            = 22,000
Actual overhead rate = (Actual Variable overhead / Actual hours)
                                           = (352,000/ 440,000)
                                          = 0.80
Standard overhead rate ( Standard Variable overhead - Standard hours)
                                          =   300,000 / 400,000
                                          = 0.75

Variable Overhead efficiency variance

a. 33,000 U

Variable overhead efficiency variance = Standard overhead rate * (Actual Hours - Standard Hours required to Produce Actual units)
                                                                              = 0.75 * (440,000 -396,000)
33,000 U
Standard hours required to Produce Actual Units = 198000*2
                = 396,000
Standard variable overhead rate        = (Standard Variable overhead / standard Hours)
                                                                         = 300,000/ 400,000
                                                                         = 0.75
Standard variable overhead = Total standard overheads - Fixed overhead)
                                                          = 900,000 - 600,000
                                                         = 300,000
Standard Fixed overhead = Standard outputs * Standard overhead per unit
                                                    = 300,000 * 2
                                                    = 600,000